OpenStudy (anonymous):

15/3 in simplest form

OpenStudy (anonymous):

5/1

OpenStudy (anonymous):

I got 5 1 over 3

OpenStudy (anonymous):

no its just 5

OpenStudy (anonymous):

ok thxs @FriedRice

OpenStudy (anonymous):

7= 21/?

OpenStudy (igreen):

21/3 = 7

OpenStudy (igreen):

5/1 = 5

OpenStudy (igreen):

So it doesn't really matter if it's 5/1 or 5..

OpenStudy (igreen):

7/21 DOES NOT = 7

OpenStudy (anonymous):

wtf are you talking about green?

OpenStudy (anonymous):

thxs

OpenStudy (igreen):

7= 21/? is the question. You said 21/7 is the answer.

OpenStudy (igreen):

But the answer is 21/3 because 3 * 7 = 21

OpenStudy (anonymous):

you you said that

OpenStudy (anonymous):

omg your such a troll

OpenStudy (anonymous):

lol

OpenStudy (igreen):

What..?

OpenStudy (anonymous):

you get the same answer

OpenStudy (igreen):

No -.- You said the answer was 7/21. I SAID IT WAS 21/3

OpenStudy (igreen):

7/21 and 21/3 ARE NOT THE SAME

OpenStudy (anonymous):

well i was talking about what you were saying in your own commnets

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

but your correct

OpenStudy (igreen):

:l

OpenStudy (anonymous):

:D